the weight of one atom of uranium is 238 amu.its actual weight is  ....gm

Grade:10

4 Answers

User
5 years ago
1amu=1.67×10^-24 so mass of one atom in grams=238×1.67×10^-241 mole have 6.022×10^23 atoms so actual weight =6.022×10^23×238×1.67×10^-24=239.35
Anjali Agrawal
5 years ago
no of molecules Mole=no of molecules ÷NAMass of uranium in gram)=238÷6.02*10^23 17÷43*10^-233.95*10^-22 *NA = mole
Anjali kumari
5 years ago
As we know 1amu =1.66×10^-24 g.So the weight of 238 amu of Uranium=238×1.66×10^-24g =3.95×10^-22g.The above answer have wrong method.
Anjali
4 years ago
1 atom of uranium=238 amu
                      1 amu=1.67 *10-24 kg
                238 amu=238*1.67*10-24/1000 g
                                 =397.46*10-24*10-3 g
                                        =3.97*10-22 g